OpenStudy (anonymous):

A community garden group is giving boxes of plants to local gardeners. The people who assembled the boxes randomly picked items from 9 kinds of vegetable plants, 5 kinds of trees, 4 types of flowers, and 3 types of herbs. Find the probability that Greg will pick a box with an asparagus plant, a white ash tree, a daylily, and a cilantro herb plant

OpenStudy (studygurl14):

\[\frac{1}{9} \times \frac{1}{5} \times \frac{1}{4} \times \frac{1}{3}=answer\]

OpenStudy (anonymous):

thx